OpenStudy (anonymous):

Two functions, f(x) = 2x2 + 3 and g(x) = x + 1. f[g(x)] = 2x2 + 4x + 5 Select one: True False

OpenStudy (nubeer):

True !!

OpenStudy (haseeb96):

f(g(x)) = f(x+1) now put x+1 in the place of x in the f(x) every place =2(x+1) ^2 +3 =2 (x^2 +2x +1) +3 = 2x^2 +4x +2 +3 = 2x^2 +4x+5 now the answer is TRUE it is same as of yurs